Calculate the area of ​​the region R in the xy plane enclosed in the circumference

x2 + y2= 4, to the right of line x = 1 and above line x =√3y.


1
Expert's answer
2021-06-15T09:27:33-0400

"x=rcos\theta,y=rsin\theta"

"r=2"


For upper limit:

"2cos\theta=1"

"\theta=\pi/3"

For lower limit:

"cos\theta=\sqrt{3}sin\theta"

"\theta=\pi/6"


"S=1/2\int r^2d\theta=1/2\int^{\pi/3}_{\pi/6}4d\theta=2(\pi/3-\pi/6)=\pi/3"